# Investigating the correlation between ZTF TDEs and IceCube high-energy neutrinos

### v1

@@ Abstract
The main contributors of the IceCube diffuse neutrino flux remain unclear. Tidal disruption events (TDEs) have been proposed as potential emitters of the high-energy neutrinos detected by IceCube. Therefore, investigating the correlation between the TDE population and IceCube neutrinos could help us better understand whether the TDE population could be potential high-energy neutrino emitters. In this paper, we perform a systematic search for TDEs that are associated with neutrinos in a sample including 143 IceCube neutrino alert events and 61 TDEs classified by the Zwicky Transient Facility (ZTF) - Bright Transient Survey (BTS). Furthermore, considering that the TDEs/TDE candidates reported as potential IceCube neutrino emitters are all accompanied by infrared (IR) observations, we further select the TDEs with IR observations from these 61 TDEs as a subsample to examine the correlation with neutrinos. Based on the Wide-field Infrared Survey Explorer (WISE) mission database, seven TDEs are identified as having IR observations. Due to good spatial localization is crucial for association analysis, we employ two methods to handle alert events with large error radii in our sample. Then we employ three Monte Carlo simulation methods to investigate the correlation between TDE sample/subsample and IceCube neutrinos. Finally, after considering spatial and temporal criteria, seven TDEs with IR flares show the most significant correlation at a 2.43 confidence level. If we tentatively further take the time delay factor into account in the weighting scheme, the correlation enhances to 2.54 confidence level.
@@ 1 Introduction
The detection of the TeV-PeV diffuse astrophysical neutrino flux by IceCube (Aartsen et al., 2013a, 2014a, 2015, 2016, 2020a; Abbasi et al., 2021a, 2022a) has unveiled new possibilities for exploration in both astrophysics (Kistler & Beacom, 2006; Beacom & Kistler, 2007; Murase et al., 2011; Murase & Ioka, 2013; Murase et al., 2013; Ahlers & Halzen, 2014; Tamborra et al., 2014; Murase et al., 2014; Bechtol et al., 2017; Kistler & Laha, 2018; Bartos et al., 2017; Sudoh et al., 2018; Bartos et al., 2019; Bustamante & Ahlers, 2019; Hovatta et al., 2021; Ackermann et al., 2022) and particle physics (Gonzalez-Garcia et al., 2005; Ioka & Murase, 2014; Ng & Beacom, 2014; Aartsen et al., 2017a; Bustamante & Connolly, 2019; Zhou & Beacom, 2020a, b, 2022; Aartsen et al., 2021; Ackermann et al., 2022; Guo et al., 2023; Plestid & Zhou, 2024; Lü et al., 2024). Despite a significant number of studies have been conducted to identify the sources of these astrophysical neutrinos. (Abbasi et al., 2011; Aartsen et al., 2013b, 2014b; Adrian-Martinez et al., 2016; Aartsen et al., 2017b, c, 2019a, 2020b, 2019b, 2020c; Lu et al., 2024; Kouch et al., 2024). It remains unclear which astrophysical sources are the principal origins of these neutrinos, although the nearby Seyfert galaxy NGC 1068 has been identified as a neutrino source at the 4.2 confidence level (Abbasi et al., 2022b).
@@
Among various astrophysical sources, tidal disruption events (TDEs) have been proposed as potential emitters of high-energy neutrinos. TDE is a class of bright transients, and the general picture of TDE involves a star being tidally disrupted when it approaches sufficiently close to a supermassive black hole (SMBH). A portion of the disrupted stellar debris is subsequently accreted by the SMBH, resulting in the production of an electromagnetic flare that occurs over timescales ranging from months to years (Rees, 1988). Numerous models have been proposed to explain the origin of the nonthermal electromagnetic and neutrino emissions from TDEs, including relativistic jets (Wang et al., 2011; Wang & Liu, 2016; Dai & Fang, 2017; Senno et al., 2017), accretion disks (Hayasaki & Yamazaki, 2019), wide-angle outflows/hidden winds (Fang et al., 2020), and tidal stream interactions (Dai et al., 2015; Hayasaki & Yamazaki, 2019).
@@
Among the TDEs/TDE candidates that have been reported as potential IceCube neutrino emitters, three notable events include: one identified TDE AT2019dsg (Stein et al., 2021), and two TDE candidates, AT2019fdr (Reusch et al., 2022) and AT2019aalc (van Velzen et al., 2024), which correspond to the possibly associated alert events IC191001A, IC200530A, and IC19119A, respectively. In these three events, AT2019dsg has been classified as a member of TDE by the Zwicky Transient Facility (ZTF). The long-lived nonthermal emission detected in AT2019dsg (Cendes et al., 2021; Stein et al., 2021) suggests that mildly relativistic outflows may provide an ideal environment for neutrino production. In contrast, the classification of AT2019fdr as a TDE is somewhat uncertain, as the flare occurred in a known active galactic nucleus (AGN), complicating its origin due to the presence of an accretion disk (e.g., Trakhtenbrot et al., 2019; Frederick et al., 2021; Cendes et al., 2021). Similarly, AT2019aalc is also posited to be a flare originating from another known AGN (van Velzen et al., 2024; Veres et al., 2024). These three events share several comparable characteristics, for instance, they have bright optical luminosities, while accompanying a delayed emission in infrared (IR) observation. Further, all of these events have a time delay for the neutrino detection, which has been explained by some models (Liu et al., 2020; Murase et al., 2020; Winter & Lunardini, 2021; Hayasaki, 2021; Wu et al., 2022; Winter & Lunardini, 2023; Mukhopadhyay et al., 2024).
@@
As more and more TDEs/TDE candidates are reported as potential IceCube high-energy neutrino emitters, investigating the correlation between the TDE population and high-energy neutrinos is helpful for understanding whether the TDE population could indeed be contributors to the IceCube neutrino flux. In this work, we investigate the hypothesis that the TDE population is correlated with IceCube alert events111https://gcn.gsfc.nasa.gov/amon_icecube_gold_bronze_events.html. We utilize the TDEs classified by ZTF - Bright Transient Survey (BTS)222https://sites.astro.caltech.edu/ztf/bts/bts.php (Perley et al., 2020), which is a public catalog with a strict pipeline to conduct a classification for transient sources, and a magnitude-limited ( in either the or filter) survey for extragalactic transients in the ZTF public stream (Perley et al., 2020). We systematically search for spatial association events between TDEs of the ZTF-BTS catalog and IceCube alert events, while allowing a time delay between TDEs and the alert events. Since IR echoes is detected from all the TDEs/TDE candidates that are reported as potential IceCube neutrino emitters, we also systemically search for the IR emission at the coordinates of ZTF-BTS TDEs using the Wide-field Infrared Survey Explorer (WISE) mission database (Wright et al., 2010). Thus, we further adopt TDEs with IR observations as a subsample to investigate the correlation with IceCube high-energy neutrinos.
@@ 2.1 TDE sample
The ZTF conducts a two-day cadence survey of the visible northern sky () with newly found transient candidates announced via public alerts. To complement the photometric survey, the ZTF-BTS undertakes an extensive spectroscopic campaign. Since 2018 April 1, ZTF-BTS has aimed to spectroscopically classify all extragalactic transients with a magnitude of m 18.5 mag (and discoveries with m 19 mag) in the g or r filters, with the classifications made publicly available (Fremling et al., 2020). The survey contains thousands of objects and is updated nearly daily. This catalog offers a comprehensive and purely magnitude-limited sample of extragalactic transients, suitable for detailed statistical and demographic analysis. The ZTF-BTS catalog sets a series of stringent quality criteria to classify transients, such as magnitude limit, observation times and Galactic extinction limit (see Perley et al. (2020) for details). Therefore, this catalog could provide several TDEs that have been reliably identified for performing this correlation analysis. We select TDEs according to the classifications reported by ZTF-BTS, including 61 TDEs333We note that AT2019fdr is classified as a supernova and AT2019aalc is not included in the ZTF-BTS catalog., as shown in Fig 2.
@@
Then we search for the IR emission from coordinates of 61 ZTF-BTS TDEs using the Near-Earth Object WISE Reactivation database444https://irsa.ipac.caltech.edu/cgi-bin/Gator/nph-dd?catalog=neowiser_p1bs_psd (NEOWISE-R; Mainzer et al., 2014; NEOWISE Team, 2020). NEOWISE-R database is based on an imaging survey centered at 3.4 and 4.6 m (labeled W1 and W2). The photometric measurements are obtained from the fitting of the point-spread function (PSF), where the PSFs are estimated from observations involving tens of thousands of stars. The NEOWISE-R magnitudes are measured through PSF fitting for individual exposures, which have been utilized to examine the IR flares for astrophysical sources (e.g., Jiang et al., 2021).
@@
The obtained single-exposure data are initially filtered according to the quality flags indicated in the catalogs. We exclude the bad data points with poor-quality frames (qifact 1), charged particle hits (saasep 5), scattered moonlight (moonmasked = 1), and artifacts (ccflags 0). Additionally, data is abandoned when multiple PSF components (nb 1 and na 0) are present during photometry, particularly when the source is fitted concurrently with other nearby detections or when a single detection is split into two components during the fitting process. To examine the variability of IR emission at the positions of TDEs, we estimate the magnitudes by adopting the median values of the data points. Then we qualify the significance of variability in IR emission at the position of each TDE and take uncertainties into consideration, as given by Jiang et al. (2021),
@@
where, and are the maximum and minimum magnitude for -th band, respectively. is the error of and is the error of . To identify the IR flare, we impose a requirement that the IR valid variability, denoted as or , must exceed 5. Meanwhile, the IR flares should only occur after the discovery dates of TDEs, where the discovery dates of TDEs are obtained from Transient Name Server (TNS)555https://www.wis-tns.org/search.
Moreover, we note that the discovery dates of five TDEs occur after the final data point of the NEOWISE-R database, namely AT2023wdb, AT2024bgz, AT2024pvu, AT2024tvd and AT2024yqo. For these five TDEs, we search for their information from Extragalactic Database (NASA/IPAC Extragalactic Database (2019), NED), and find no report indicating that they accompany IR flares. Finally, there are seven ZTF-BTS TDEs with IR flares, as listed in Table 1. The light curves of these seven TDEs are presented in Fig 1, where red and blue circles are the W1 and W2 band data, respectively. The red dashed lines are the discovery dates of TDEs.
TNSID | Discovery datea | MJD | ||
AT2019azh | 19/02/22 | 58535 | 5.15 | 3.48 |
AT2019dsg | 19/04/09 | 58582 | 18.59 | 16.40 |
AT2019qiz | 19/09/19 | 58745 | 38.87 | 41.78 |
AT2020nov | 20/06/27 | 59027 | 12.95 | 9.30 |
AT2022dyt | 22/02/26 | 59635 | 7.96 | 7.37 |
AT2022upj | 22/08/31 | 59822 | 30.36 | 16.17 |
AT2023cvb | 23/03/06 | 60009 | 7.76 | 6.77 |
@@
b The significance of the IR variability in the W1 and W2 bands.
@@ 2.2 Neutrino sample
The IceCube Neutrino Observatory has observed high-energy astrophysical neutrinos exceeding the atmospheric background in both cascade and track data. Since 2016, real-time alerts for individual high-energy events (100 TeV) have been released to the multimessenger observational community through platforms, for example, the Global Cycling Network (GCN)666https://gcn.nasa.gov/circulars?query=IceCube&startDate=&endDate=&sort=circularID&limit=100. These alert events concentrate on track-like neutrino candidates, which provide more precise angular localizations compared to cascade events. Since June 2019, all alert events have been assigned a signalness value and categorized as either gold or bronze based on their probability of astrophysical origin, specifically larger than 50 and 30, respectively. The 90 containment error radius (777Noting that the only contains statistical error as described in https://gcn.gsfc.nasa.gov/amon_icecube_gold_bronze_events.html) for these alert events are publicly available. For each alert event, an updated report would be provided after detecting a more accurate position. The latest reported information for each alert event is adopted in our analysis, while we note that one alert event888https://gcn.nasa.gov/circulars/37794?query=IceCube&startDate=&endDate=&sort=circularID&limit=100 (i.e. IC241016) is found to clearly not be of astrophysical origin so that we remove it from our sample. We also note that Abbasi et al. (2023a) has provided updated for neutrinos between 2011 and 2020. However, compared to the catalog covering the period from 2011 to 2020, our primary focus is on the neutrinos published after 2019, as this period suitably matches the ZTF-BTS catalog survey timeline i.e. allowing for alert events occur after the TDEs of ZTF-BTS catalog. Therefore, for our analysis, we adopt the list of all the bronze and gold alert events after 2019 as our sample, leaving at least a 30 probability of an astrophysical origin for these alert events.

### v2

@@ Abstract
The main contributors of the IceCube diffuse neutrino flux remain unclear. Tidal disruption events (TDEs) have been proposed as potential emitters of the high-energy neutrinos detected by IceCube. Therefore, investigating the correlation between the TDE population and IceCube neutrinos could help us better understand whether the TDE population could be potential high-energy neutrino emitters. In this paper, we perform a systematic search for TDEs that are associated with neutrinos in a sample including 143 IceCube neutrino alert events and 56 TDEs classified by the Zwicky Transient Facility (ZTF) - Bright Transient Survey (BTS). Furthermore, considering that the TDEs/TDE candidates reported as potential IceCube neutrino emitters are all accompanied by infrared (IR) observations, we further select the TDEs with IR observations from these 56 TDEs as a subsample to examine the correlation with neutrinos. Based on the Wide-field Infrared Survey Explorer (WISE) mission database, eight TDEs are identified as having IR observations. Since good spatial localization is crucial for association analysis, we employ two methods to handle alert events with large error radii in our sample. Then we employ Monte Carlo simulations to quantify the correlation between the TDE sample/subsample and IceCube neutrinos. Finally, after considering spatial and temporal criteria, the eight TDEs with IR echoes show the most significant correlation at a 2.37 confidence level. If we tentatively further take the time delay factor into account in the weighting scheme, the correlation enhances to 2.65 confidence level.
@@ 1 Introduction
The detection of the TeV-PeV diffuse astrophysical neutrino flux by IceCube (Aartsen et al., 2013a, 2014a, 2015, 2016, 2020a; Abbasi et al., 2021a, 2022a) has unveiled new possibilities for exploration in both astrophysics (Kistler & Beacom, 2006; Beacom & Kistler, 2007; Murase et al., 2011; Murase & Ioka, 2013; Murase et al., 2013; Ahlers & Halzen, 2014; Tamborra et al., 2014; Murase et al., 2014; Bechtol et al., 2017; Kistler & Laha, 2018; Bartos et al., 2017; Sudoh et al., 2018; Bartos et al., 2019; Bustamante & Ahlers, 2019; Hovatta et al., 2021; Ackermann et al., 2022) and particle physics (Gonzalez-Garcia et al., 2005; Ioka & Murase, 2014; Ng & Beacom, 2014; Aartsen et al., 2017a; Bustamante & Connolly, 2019; Zhou & Beacom, 2020a, b, 2022; Aartsen et al., 2021; Ackermann et al., 2022; Guo et al., 2023; Plestid & Zhou, 2024; Lü et al., 2024). Despite a significant number of studies have been conducted to identify the sources of these astrophysical neutrinos (Abbasi et al., 2011; Aartsen et al., 2013b, 2014b; Adrian-Martinez et al., 2016; Aartsen et al., 2017b, c, 2019a, 2020b, 2019b, 2020c; Lu et al., 2024; Kouch et al., 2024), it remains unclear which astrophysical sources are the primary origins of these neutrinos, and only a very small number of sources (e.g., the Seyfert galaxy NGC 1068) have been identified as neutrino sources (Abbasi et al., 2022b).
@@
Among various astrophysical sources, tidal disruption events (TDEs) have been proposed as potential emitters of high-energy neutrinos. TDE is a class of bright transients, and the general picture of TDE involves a star being tidally disrupted when it approaches sufficiently close to a supermassive black hole (SMBH). A portion of the disrupted stellar debris is subsequently accreted by the SMBH, resulting in the production of an electromagnetic flare that occurs over timescales ranging from months to years (Rees, 1988). Numerous models have been proposed to explain the origin of the nonthermal electromagnetic and neutrino emissions from TDEs, including relativistic jets (Wang et al., 2011; Wang & Liu, 2016; Dai & Fang, 2017; Senno et al., 2017; Lunardini & Winter, 2017; Liu et al., 2020; Hayasaki, 2021), accretion disks (Hayasaki & Yamazaki, 2019), wide-angle outflows/hidden winds (Fang et al., 2020), and tidal stream interactions (Dai et al., 2015; Hayasaki & Yamazaki, 2019).
@@
Among the TDEs/TDE candidates that have been reported as potential IceCube neutrino emitters, three notable events include: one identified TDE AT2019dsg (Stein et al., 2021), and two TDE candidates, AT2019fdr (Reusch et al., 2022) and AT2019aalc (van Velzen et al., 2024), which correspond to the possibly associated alert events IC191001A, IC200530A, and IC19119A, respectively. In these three events, AT2019dsg has been classified as a member of TDE by the Zwicky Transient Facility (ZTF). The long-lived nonthermal emission detected in AT2019dsg (Cendes et al., 2021; Stein et al., 2021) suggests that mildly relativistic outflows may provide an ideal environment for neutrino production. In contrast, the classification of AT2019fdr as a TDE is somewhat uncertain, as the flare occurred in a known active galactic nucleus (AGN), complicating its origin due to the presence of an accretion disk (e.g., Trakhtenbrot et al., 2019; Frederick et al., 2021; Cendes et al., 2021). Similarly, AT2019aalc is also posited to be a flare originating from another known AGN (van Velzen et al., 2024; Veres et al., 2024). These three events share several comparable characteristics; for instance, they have bright optical luminosities, accompanying infrared (IR) echoes, i.e., delayed IR rising after the discovery date of TDE. Further, all of these events have a time delay for the neutrino detection, which has been explained by some models, for instance, protons generated by the delayed choked jet (Mukhopadhyay et al., 2024), the slow acceleration of protons by the plasma/magnetic reconnection (Murase et al., 2020), an increase of the neutrino production efficiency over time assuming that the radius of neutrino production evolves similarly to that of the black body (Winter & Lunardini, 2021), and the protons accelerated by the outflow/jet interacting with protons/photons after traveling to the cloud/dust region (Wu et al., 2022; Winter & Lunardini, 2023).
@@
As more and more TDEs/TDE candidates are reported as potential IceCube high-energy neutrino emitters, investigating the correlation between the TDE population and high-energy neutrinos is helpful for understanding whether the TDE population could indeed be contributors to the IceCube neutrino flux. In this work, we investigate the hypothesis that the TDE population is correlated with IceCube alert events111https://gcn.gsfc.nasa.gov/amon_icecube_gold_bronze_events.html. We utilize the TDEs classified by ZTF - Bright Transient Survey (BTS)222https://sites.astro.caltech.edu/ztf/bts/bts.php (Perley et al., 2020), which is a public catalog with a strict pipeline to conduct a classification for transient sources, and a magnitude-limited ( in either the or filter) survey for extragalactic transients in the ZTF public stream (Perley et al., 2020). We systematically search for spatial association events between TDEs of the ZTF-BTS catalog and IceCube alert events, while allowing a time delay between TDEs and the alert events. Since IR echoes are detected from all the TDEs/TDE candidates that are reported as potential IceCube neutrino emitters, we also systematically search for the IR emission at the coordinates of ZTF-BTS TDEs using the Wide-field Infrared Survey Explorer (WISE) mission database (Wright et al., 2010). Thus, we further adopt TDEs with IR observations as a subsample to investigate the correlation with IceCube high-energy neutrinos.
@@ 2.1 TDE sample
Since 2018, the ZTF has conducted a three-day cadence survey of the visible northern sky (), announcing newly found transient candidates via public alerts (Masci et al., 2018; Graham et al., 2019; Bellm et al., 2019; Fremling et al., 2020). In December 2020, the ZTF public survey has transitioned to Phase-II operations, with the survey shifting to a two-day cadence. To complement the photometric survey, the ZTF-BTS undertakes an extensive spectroscopic campaign, aiming to spectroscopically classify all extragalactic transients with a peak magnitude of mag in the or filters (transients down to 19.0 mag are also included when spectroscopic resources are available), with the classifications made publicly available (Fremling et al., 2020). The survey contains over 10,000 objects and is updated nearly daily2. This catalog offers a comprehensive and purely magnitude-limited sample of extragalactic transients, suitable for detailed statistical and demographic analysis. The ZTF-BTS catalog sets a series of stringent quality criteria to classify transients, such as magnitude limit, observation times and Galactic extinction limit (see Perley et al. (2020) for details). Therefore, this source catalog can provide some TDEs/candidates with relatively reliable classification as our sample for performing the correlation analysis. We select TDEs according to the classifications reported by ZTF-BTS, including 61 TDEs up to November 2024333Note that AT2019fdr is classified as a supernova and AT2019aalc is not included in the ZTF-BTS catalog..
@@
Then we search for the IR echoes from coordinates of ZTF-BTS TDEs using the full Near-Earth Object WISE Reactivation database444https://irsa.ipac.caltech.edu/cgi-bin/Gator/nph-dd?catalog=neowiser_p1bs_psd (NEOWISE-R; Mainzer et al., 2014; NEOWISE Team, 2020). The full NEOWISE-R database is up to July 2024. To ensure that the IR echoes of TDEs can be included in the NEOWISE-R database, we limit our TDE sample to before May 2024, due to the fact that the IR echoes of TDEs often have significant time delay. After the time cut, there are 56 TDEs remaining in our sample, as shown in Fig. 1. The NEOWISE-R database is based on an imaging survey centered at 3.4 and 4.6 m (labeled W1 and W2). The photometric measurements are obtained from the fitting of the point-spread function (PSF), where the PSFs are estimated from observations involving tens of thousands of stars. The NEOWISE-R magnitudes are measured through PSF fitting for individual exposures, which have been utilized to examine the IR flares for astrophysical sources (e.g., Jiang et al., 2021).
@@
The obtained single-exposure data are initially filtered according to the quality flags indicated in the catalogs. We exclude the bad data points with poor-quality frames (qi_fact 1), charged particle hits (saa_sep 5), scattered moonlight (moon_masked 1), and artifacts (cc_flags 0). Additionally, data is abandoned when multiple PSF components (nb 1 and na 0) are present during photometry, particularly when the source is fitted concurrently with other nearby detections or when a single detection is split into two components during the fitting process. To examine the variability of IR emission at the positions of TDEs, we estimate the magnitudes by adopting the median values of the data points. Then we qualify the significance of variability in IR emission at the position of each TDE and take uncertainties into consideration, as given by Jiang et al. (2021),
@@
where, and are the maximum and minimum magnitude for -th band, respectively. is the error of and is the error of . To identify the IR variability, we impose a requirement that the IR variability, denoted as or , must exceed 5. We used the following criteria to further select TDEs with IR echoes from these TDEs with IR variability. 1) The IR magnitude presents brightening compared to the median value of the quiescent level before the discovery date (Jiang et al., 2021). 2) The brightest magnitude should occur after the discovery dates of TDEs that are obtained from the Transient Name Server (TNS)555https://www.wis-tns.org/search. Finally, there are 8 ZTF-BTS TDEs with IR echoes, as listed in Table 1. The light curves of these eight TDEs are presented in Fig 2, where red and blue circles are the W1 and W2 band data, respectively. The red dashed lines are the discovery dates of TDEs.
TNSID | Discovery datea | MJD | log | log | |||||
(days) | (mag) | (mag) | erg s-1 | erg s-1 | |||||
AT2019azh | 19/02/22 | 58535 | 5.15 | 3.48 | 0.15 | 0.20 | 0.022 | 42.04 | 41.85 |
AT2019dsg | 19/04/09 | 58582 | 18.59 | 16.40 | 0.86 | 1.51 | 0.0512 | 42.60 | 42.69 |
AT2019qiz | 19/09/19 | 58745 | 38.87 | 41.78 | 1.32 | 2.38 | 0.0151 | 43.21 | 43.41 |
AT2020nov | 20/06/27 | 59027 | 12.95 | 9.30 | 0.56 | 0.76 | 0.084 | 42.31 | 42.28 |
AT2020afhd | 24/01/13 | 60322 | 5.76 | 3.96 | 0.20 | 0.18 | 0.027 | 41.92 | 41.58 |
AT2022dyt | 22/02/26 | 59635 | 7.96 | 7.37 | 0.46 | 0.79 | 0.072 | 42.14 | 42.18 |
AT2022upj | 22/08/31 | 59822 | 30.36 | 16.17 | 1.66 | 2.29 | 0.054 | 42.96 | 42.96 |
AT2023cvb | 23/03/06 | 60009 | 7.76 | 6.77 | 0.63 | 1.05 | 0.071 | 42.27 | 42.33 |
@@
b The significance of the IR variability between the maximum and minimum magnitudes in the W1 and W2 bands.
c The magnitude difference between the brightest magnitude and the median value of quiescent level.
d The luminosity difference between the brightest magnitude and the median value of quiescent level.
@@ 2.2 Neutrino sample
The IceCube Neutrino Observatory has observed high-energy astrophysical neutrinos exceeding the atmospheric background in both cascade and track data. Since 2016, real-time alerts for individual high-energy events (100 TeV) have been released to the multimessenger observational community through platforms, for example, the Global Cycling Network (GCN)666https://gcn.nasa.gov/missions/icecube. These alert events concentrate on track-like neutrino candidates, which provide more precise angular localizations compared to cascade events. Since June 2019, all alert events have been assigned a signalness value and categorized as either gold or bronze based on their probability of astrophysical origin, specifically larger than 50 and 30, respectively. The 90 containment error radii (, statistical error only) for these alert events are publicly available. For each alert event, an updated report would be provided after detecting a more accurate position. The latest reported information for each alert event is adopted in our analysis, while we note that one alert event (i.e., IC241016777https://gcn.nasa.gov/circulars/37794) is found to clearly not be of astrophysical origin so we remove it from our sample. We also note that Abbasi et al. (2023a) has provided updated for neutrinos between 2011 and 2020. However, compared to the catalog covering the period from 2011 to 2020, our primary focus is on the neutrinos published after 2019, as this period suitably matches the ZTF-BTS catalog survey timeline i.e. allowing for alert events occur after the TDEs of ZTF-BTS catalog. Therefore, for our analysis, we adopt the list of all the bronze and gold alert events after 2019 as our sample, leaving at least a 30 probability of an astrophysical origin for these alert events.